Help with this one too?
This linear systems by substitution, i think this is a no solution problem though , i have tried everything and i can not get it . I have literally been on this question for 3 months now.....
i may just have a big brain fart or something but here it is.
3x+5y=12
x+3y=6
: Help with this one too?
This linear systems by substitution, i think this is a no solution problem though , i have tried everything and i can not get it . I have literally been on this question for 3 months now.....
i may just have a big brain fart or something but here it is.
3x+5y=12
x+3y=6

Answer by scott8148(2719)

from 2nd eqn __ x=6-3y

substituting into 1st eqn __ 3(6-3y)+5y=12 __ 18-9y+5y=12 __ -4y=-6 __ y=3/2

substituting __ x+3(3/2)=6 __ x+9/2=6 __ x=3/2

so x and y are equal __ maybe that was causing your difficulty